OpenStudy (anonymous):

Substitution Method What is the solution to the system of equations? 4x + y = 2 x + y = 3 (_,_)

OpenStudy (anonymous):

@aaronq @Abhisar @adrynicoleb @AnswerMyQuestions @animal_lover36 @ASAAD123 @blurbendy @boricua5 @Blu-Girl @Brostep0s @briananicolenelson @Bigtim10 @campbell_st @cwrw238 @Compassionate @confluxepic @Conqueror

OpenStudy (confluxepic):

@TheSmartOne

OpenStudy (confluxepic):

Do you know how to graph. @pumkinspicelatte!

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

The answer is where the lines touch: https://www.desmos.com/calculator/5iwcsaui9l

OpenStudy (anonymous):

so its (1,3)?

OpenStudy (anonymous):

No. Where do the lines touch?

OpenStudy (anonymous):

https://www.desmos.com/calculator/medhe7n1o8

OpenStudy (anonymous):

(1.5,3.5)?

OpenStudy (anonymous):

No....

OpenStudy (anonymous):

Click at the place where the lines touch, and it will tell you.

OpenStudy (anonymous):

ok I did it but that's way too long of a number that cant be right

OpenStudy (anonymous):

It is the answer.

OpenStudy (anonymous):

start by solving for x or y using equation 2 x+y=3 y=3-x now plug that into your top equation to solve for x 4x + (3-x)=2 4x +3 -x=2 3x+3=2 3x=-1 x=-1/3 now solve for y 4(-1/3) + y=2 -4/3+y=2 y=10/3 (-1/3, 10/3)